A half-liter of water, initially at 30◦C, is cooled by removing 63 kJ of heat energy. What is the final temperature of the water? in C a.79.1 b.45.1 c.0 d.15 e.12
Answer #1

Option C

Mass of water = 500 g ( density of water= 1 g / ml)

Initial temperature of water =30^{o }C

Final temperature of the system= t

Specific heat of water = 4.184j/g. ^{o}C

Energy removed = 63 kj

= 63000 joules

Heat energy content removed, Q = mass of water * specific heat of water * temperature difference

63000 j = 500 g * 4.184 j /goC * dT

dT = 63000/ 500 *4.184

= 12.6/ 4.184

=30.11

Final temperature = 30 - 30.11

= 0.11oC

Best Option C